McDormand, Inc., reported a $3,500 unfavorable price variance for variable overhead and a $35,000 unfavorable price variance for fixed overhead. The flexible budget had $1,084,500 variable overhead based on 36,150 direct labor-hours; only 34,110 hours were worked. Total actual overhead was $1,829,800. The number of estimated hours for computing the fixed overhead application rate totaled 38,400 hours.  
  

Required:

a. Prepare a variable overhead analysis.
b. Prepare a fixed overhead analysis.
Answer #1

A)  price variance 3,500 U

1,084,500/ 36,150= 30            34,110* 30= 1,023,300

efficiency variance 1,084,500- 1,023,300= 61,200 F


variable OH cost variance= 61,200- 3,500= 57,700 F


B) price variance= 35,000 U

 or 803,000- 768,000= 35,000 U


1,829,800- 1,026,800= 803,000

803,000- 35,000= 768,000

768,000/ 38,400= 20

20* 36,150= 723,000

Production volume variance= 768,000- 723,000= 45,000 U

Fixed overhead cost variance= 35,000+ 45,000= 80,000 U